A billboard is on top of a building. You are 134.9 feet from the building. You measure the angle of elevation of the bottom of the billboard as 39.4 degrees and the angle of elevation of the top of the billboard as 47degrees. How tall is the billboard?
Answer by mananth(16946)

let bill board height be x
building height =y
Tan 47 = (x+y)/134.9
Tan 39.4 = y/134.9
Tan 47-Tan 39.4 = (x+y)/134.9 - y/134.9
Tan 47-Tan39.4 = x/134.9
x= 134.9(Tan 47-tan39.4)
x= 33.85 feet